11. The 4077th Mobile Army Surgical Hospital hospital is considering the purchase of a helicopter to transport critical patie I need help with question 11 part A,B, C, D, and E
Answer #1

a)

for this to be valid:

\sumP(x) =P(X=0)+P(X=1)+P(X=2)+P(X=3)+P(X=4)+P(X=5)+P(X=6) must be1

\sumP(x) =P(X=0)+P(X=1)+P(X=2)+P(X=3)+P(X=4)+P(X=5)+P(X=6)=0.15+0.2+0.34+0.19+0.06+0.05+0.01

=1

since all individual sample point probabilities are between 0 and 1 and sum of probability of all sample point is 1, this is a valid probability distribution

b)

E(X)=\sumxP(x) =0*0.15+1*0.2+2*0.34+3*0.19+4*0.06+5*0.05+6*0.01=2

c)

E(X2)=\sumx2P(x) =0^2*0.15+1^2*0.2+2^2*0.34+3^2*0.19+4^2*0.06+5^2*0.05+6^2*0.01=5.84

Var(x)=σ2 = E(x2)-(E(x))2= 1.8400
std deviation=         σ= √σ2 = 1.3565

d)

P(X=0) =0.15

e)\

P(X>=1) =1-P(X=0)=1-0.15 =0.85